OpenStudy (anonymous):

Write the solution in lowest terms. 1/3 - (-3/4) =

OpenStudy (anonymous):

This is the same as (1/3) + (3/4) = 13/12

OpenStudy (anonymous):

1/3 - (-3/4) is the same as 1/3 + 3/4 Find the common denominator. 4/12 + 9/12 = 13/12

OpenStudy (anonymous):

Find the common denominator of 3 and 4, which is 12. 1/3+3/4= (4+9)/12 = 13/12